# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

### v4

@@ V-D2 Task Collaboration and Interference
We observe several key findings. First, removing any task generally leads to lower average performance of human-centric tasks. Particularly, removing the 2D Pose task significantly decreased performance in detection (-16.8 AP), mesh recovery (increased error by 18.6 MPVPE), and 3D pose estimation (increased error by 16.7 MPJPE). The 2D keypoint information in 2D pose estimation is highly important to the 3D keypoint for 3D pose estimation and can help the model perceive the boundary of pedestrians for pedestrian detection, making the 2D pose estimation task an important complement to training a human-centric foundation model. Second, tasks such as the performances of pedestrian detection, mesh recovery, and 3D pose estimation are more sensitive to the removal of other tasks. Their performance markedly declined, indicating a high interdependence among these human-centric tasks during training. Third, skeleton-based action recognition was less affected by the absence of other tasks, which may stem from its unique input type (skeleton points) compared to the image-based input of other tasks. It is also worth noting that the absence of certain tasks resulted in performance gains for human parsing, indicating potential task conflicts with other human-centric tasks that require further investigation.
TABLE XII: Alabtion of collaboration and interference between tasks. In green and red are the gaps of at least ±$\pm 0.5$ points.
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-Sub 3DPW 3DPW
mIoU AP AP mA B@4 Acc 100-MPVPE 100-MPJPE
baseline 56.8 25.5 81.8 77.9 28.0 93.2 -0.3 13.2 -
[…]
@@ V-D3 Attention Mask Designs in the Decoder
The receptive field of the attention module in the decoder is crucial for the translation and de-tokenization of features in the unified model. Therefore, we investigate the effects of different attention masks as illustrated in Fig. 5: (a) our Hulk model adopts task-specialized attention masks for different tasks, i.e., for pose estimation and mesh recovery tasks, we employ full attention; for parsing, detection, attribute, and skeleton-based tasks, we use diagonal attention; for caption tasks, we apply a combination of causal and diagonal attention. We also experimented with (b) all tasks using full attention and (c) all tasks using diagonal attention. The results, shown in Table XIII, indicate that employing different attention interaction methods for different tasks can improve the performances, i.e., +0.4% and +1.0% average performance gain on full attention and diagonal attention, respectively.
Figure 5: Given $N$ encoded tokens $\mathcal{E}(\mathbf{p})$ with $N^{\prime}$ modality indicators $\mathcal{I}_{m^{\prime}}$, we proposed various attention mask in the decoder $\mathcal{D}$ to facilitate diverse tasks. Left: Full attention enables comprehensive interaction between encoder tokens and modality indicators. Mid: Diagonal attention mitigates potential distortions in encoded tokens due to randomly initialized modality indicators. Right: A combination of the causal and diagonal masks for image captioning.
TABLE XIII: Comparison of different attention masks.
@@
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-Sub 3DPW 3DPW
@@
V-D4 Scaling up data size
To assess the impact of the diverse and sufficient human-centric data, we train Hulk using a longer 60000-iteration schedule on the ablation datasets comprising 1.7M training samples, and full datasets comprising 30M training samples, respectively. Results in Table XIV indicate that incorporating more human-centric data brings considerable performance improvement across all tasks, with notable enhancements in 2D pose estimation (+2.3% AP), pedestrian detection (+1.6% AP), image caption (+2.1 B@4), mesh recovery (-26.0 MPVPE) and 3D pose estimation (-24.5 MPJPE). Slight improvements are also noted in attribute recognition and skeleton-based action recognition. We attribute these modest gains to the lower quality of data and the significant domain gap in the extended datasets. For instance, in attribute recognition, we include LU-Person [6] with its pseudo labels that offers less informative supervision compared to other training datasets. Similarly, the Kinetics-400 dataset, which has a substantial domain gap from NTU60, influences the results for the skeleton-based action recognition.
TABLE XIV: Performance of Hulk on diverse human-centric tasks when scaling up data size. Experiments are both trained with a FULL 60,000-iteration schedule, showing that learn from additional data and achieve better performance on average.
@@ V-D4 Scaling up data size
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-Sub 3DPW 3DPW
@@
TABLE XV: Comparison of different initialization parameters.
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-Sub 3DPW 3DPW
mIoU AP AP mA B@4 Acc 100-MPVPE 100-MPJPE
(a) MAE [99] 56.8 25.5 81.8 77.9 28.0 93.2 -0.3 13.2 47.2
(b) HAP [31] 63.3 33.4 85.4 81.8 29.6 91.5 9.6 22.9 52.2
(c) PATH [30] 63.6 34.5 88.6 80.2 29.9 92.8 8.3 21.3 52.4
@@ V-D5 Different initialization weights
To examine the influence of model initialization on performance, we conduct additional experiments using pre-training parameters from the human-centric models, i.e., HAP[31] and PATH[30], and compare them with our baseline Hulk initialized by MAE [99]. The results, detailed in Table XV, indicate that initializing Hulk with parameters from human-centric pretraining methods can further enhance its accuracy. This result not only validates the robustness of Hulk but also highlight its potential for further enhancement. Moreover, it is important to note that for a fair comparison with other human-centric models, we refrain from using these specialized pretraining parameters as our default initialization method in main results, ensuring the improvement observed in Hulk are attributed to its novel architecture and learning strategy.
@@ VI Conclusion and Limitations
In this paper, we introduce Hulk, a novel generalist human-centric perceiver that marks the first integration of major 2D and 3D vision tasks, skeleton-based models, and vision-language tasks into a unified framework. Diverging from traditional models that rely on task-specific heads, Hulk adopts an innovative approach by decomposing these into two basic heads, thereby enhancing adaptability and simplifying the process. Trained on extensive datasets, Hulk not only achieves state-of-the-art performance across various tasks but does so without the need for dataset-specific fine-tuning. Beyond the success in human-centric tasks, we also shed light on the essential components for developing a human-centric foundation model. Hulk represents a significant step for human-centric foundation models, offering a versatile and effective framework that paves the way for advanced research.
@@ Acknowledgement
This work is partially supported by the National Key R&\&&D Program of China(NO.2022ZD0160101), National Natural Science Foundation of China (No.62127803), and Key R&\&&D Project of Zhejiang Province (No.2022C01056).
@@ Appendix A Full Ablation results
Appendix B Additional Architecture Details
B-A Modality indicator
B-A1 Weight sharing.
B-A2 Shape of modality indicators.
B-B Positional Embeddings
B-B1 Positional embeddings in the encoder
B-B2 Positional embeddings in the decoder
@@ B-C Objective function details
TABLE XVIII: Statistics of Training Datasets in the full setting.

### v5

@@ V-D2 Task Collaboration and Interference
We observe several key findings. First, while the removal of certain tasks can result in performance improvements in some cases, in most situations, removing one task from training will lead to an average performance decline. Among all tasks, 2D pose recognition, as a generic task, is mostly beneficial for all others. Particularly, removing the 2D Pose task significantly decreased performance in detection (-16.8 AP), mesh recovery (increased error by 18.6 MPVPE), and 3D pose estimation (increased error by 16.7 MPJPE). The 2D keypoint information in 2D pose estimation is highly important to the 3D keypoint for 3D pose estimation and can help the model perceive the boundary of pedestrians for pedestrian detection, making the 2D pose estimation task an important complement to training a human-centric foundation model. Second, tasks such as the performances of pedestrian detection, mesh recovery, and 3D pose estimation are more sensitive to the removal of other tasks. Their performance markedly declined, indicating a high interdependence among these human-centric tasks during training. Third, skeleton-based action recognition was less affected by the absence of other tasks, which may stem from its unique input type (skeleton points) compared to the image-based input of other tasks. It is also worth noting that the absence of certain tasks resulted in performance gains for human parsing, indicating potential task conflicts with other human-centric tasks that require further investigation.
Figure 5: Visualization of attribute recognition on PA-100K [101]. The missed and incorrectly identified attributes are marked in blue and red, respectively.
Figure 6: Visualization of pedestrian detection on CrowdHuman [102]. The dashed boxes indicate missed detections.
After further exploring how generic tasks help the smaller ones during Hulk training, we find that the relations between them determine their mutual help. A strong correlation between the general task and the small task amplifies the benefits of the large task, and conversely. We train Hulk with different pretraining datasets for 20,000 iterations. As shown in Table XIII involving generic tasks (2D/3D pose estimation) does not necessarily improve performance on smaller tasks, e.g., pedestrian detection, attribute recognition, and skeleton-based action recognition tasks. Specifically, adding 2D/3D pose estimation task brings $\textbf{+19.0\%}AP$$\textbf{+5.2\%}AP$ and $\textbf{+3.3\%}mA$$\textbf{+1.1\%}mA$ performance gain on CrowdHuman and RAPv2, respectively. We suggest that the information in human joint location annotations in 2D and 3D pose estimation helps Hulk to better focus on the region of pedestrians, facilitating its performance on detection and attribute recognition. In contrast, the skeleton-based action recognition task, which focuses on the semantics of human motion sequences, has a weak correlation with the joint location information learned by 2D/3D pose estimation tasks, resulting in slight performance improvements. Qualitative results in Fig. 5 and 6 also show that jointly training with generic tasks can benefit the learning on pedestrian detection and attribute recognition.
@@ V-D3 Attention Mask Designs in the Decoder
The receptive field of the attention module in the decoder is crucial for the translation and de-tokenization of features in the unified model. Therefore, we investigate the effects of different attention masks as illustrated in Fig. 7: (a) our Hulk model adopts task-specialized attention masks for different tasks, i.e., for pose estimation and mesh recovery tasks, we employ full attention; for parsing, detection, attribute, and skeleton-based tasks, we use diagonal attention; for caption tasks, we apply a combination of causal and diagonal attention. We also experimented with (b) all tasks using full attention and (c) all tasks using diagonal attention. The results, shown in Table XIV(a-c), indicate that employing different attention interaction methods for different tasks can improve the performances, i.e., +0.4% and +1.0% average performance gain on full attention and diagonal attention, respectively.
Figure 7: Given NN encoded tokens $\mathcal{E}(\mathbf{p})$ with $N^{\prime}$ modality indicators $\mathcal{I}_{m^{\prime}}$, we proposed various attention mask in the decoder $\mathcal{D}$ to facilitate diverse tasks. Left: Full attention enables comprehensive interaction between encoder tokens and modality indicators. Mid: Diagonal attention mitigates potential distortions in encoded tokens due to randomly initialized modality indicators. Right: A combination of the causal and diagonal masks for image captioning.
TABLE XIV: Comparison of different attention masks and different initialization parameters.
@@
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-XSub 3DPW 3DPW
@@
Different initialization parameters
(d) HAP [31] 63.3 33.4 85.4 81.8 29.6 91.5 9.6 22.9 52.2
(e) PATH [30] 63.6 34.5 88.6 80.2 29.9 92.8 8.3 21.3 52.4
TABLE XV: Performance of Hulk on diverse human-centric tasks when scaling up data size. Experiments are both trained with a FULL 60,000-iteration schedule, showing that learn from additional data and achieve better performance on average.
@@
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-XSub 3DPW 3DPW
@@
V-D4 Scaling up data size
To assess the impact of the diverse and sufficient human-centric data, we train Hulk using a longer 60000-iteration schedule on the ablation datasets comprising 1.7M training samples, and full datasets comprising 30M training samples, respectively. Results in Table XV indicate that incorporating more human-centric data brings considerable performance improvement across all tasks, with notable enhancements in 2D pose estimation (+2.3% AP), pedestrian detection (+1.6% AP), image caption (+2.1 B@4), mesh recovery (-26.0 MPVPE) and 3D pose estimation (-24.5 MPJPE). Slight improvements are also noted in attribute recognition and skeleton-based action recognition. We attribute these modest gains to the lower quality of data and the significant domain gap in the extended datasets. For instance, in attribute recognition, we include LU-Person [6] with its pseudo labels that offers less informative supervision compared to other training datasets. Similarly, the Kinetics-400 dataset, which has a substantial domain gap from NTU60, influences the results for the skeleton-based action recognition.
@@ V-D5 Different initialization weights
To examine the influence of model initialization on performance, we conduct additional experiments using pre-training parameters from the human-centric models, i.e., HAP[31] and PATH[30], and compare them with our baseline Hulk initialized by MAE [118]. The results, detailed in Table XIV(d-e), indicate that initializing Hulk with parameters from human-centric pretraining methods can further enhance its accuracy. This result not only validates the robustness of Hulk but also highlights its potential for further enhancement. Moreover, it is important to note that for a fair comparison with other human-centric models, we refrain from using these specialized pretraining parameters as our default initialization method in main results, ensuring the improvement observed in Hulk are attributed to its novel architecture and learning strategy.
Figure 8: Qualitative comparison of Hulk and other methods on human parsing.
Figure 9: Qualitative comparison on 2D pose estimation.
V-E Visualization
Fig. 8 and Fig. 9 show qualitative comparisons of Hulk with other SOTA methods on human parsing and 2D pose estimation, respectively. Hulk achieves more precise predictions, especially on detailed regions. More visualization can be found in the Supplementary Material.
@@ VI Conclusion and Limitations
In this paper, we introduce Hulk, a novel generalist human-centric perceiver that marks the first integration of common 2D and 3D vision tasks, skeleton-based models, and vision-language tasks into a unified framework. Diverging from traditional models that rely on task-specific heads, Hulk adopts an innovative approach by decomposing these into two basic heads, thereby enhancing adaptability and simplifying the process. Trained on extensive datasets, Hulk not only achieves state-of-the-art performance across various tasks but does so without the need for dataset-specific fine-tuning. Beyond the success in human-centric tasks, we also shed light on the essential components for developing a human-centric foundation model. Hulk represents a significant step for human-centric foundation models, offering a versatile and effective framework that paves the way for advanced research.
@@ Acknowledgement
This work was supported by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, the Research Grants Council of Hong Kong (Project No. CUHK14213224). This work was also partially supported by the National Key R&\&D Program of China(NO.2022ZD0160101), National Natural Science Foundation of China (No.62127803), and Key R&\&D Project of Zhejiang Province (No.2022C01056).
@@ Appendix A Full Ablation results
Appendix B Additional Architecture Details
B-A Modality indicator
B-A1 Weight sharing.
B-A2 Shape of modality indicators.
B-B Positional Embeddings
B-B1 Positional embeddings in the encoder
B-B2 Positional embeddings in the decoder
@@ B-C Objective function details
Appendix C Potential Negative Impacts
Appendix D Transfer to other human-centric datasets
D-A COCO-Wholebody
To evaluate the ability of Hulk on fine-grained whole-body pose estimation task, we finetune Hulk on COCO-WholeBody V1.0 dataset [145] following most of the default training and evaluation settings of mmpose [131]. Table XVIII shows the experimental results comparing Hulk with the state-of-the-art top-down methods. The results show that Hulk achieves the best whole-body pose estimation accuracy (61.2% AP and 72.1% AR) across different methods with diverse backbones. When compared using the same plain backbone, Hulk outperforms ViTPose [127] by a large margin (+5.3 AP), demonstrating the general capability of Hulk on human-centric tasks.
[…]
Method Backbone Body AP Body AR Foot AP Foot AR Face AP Face AR Hand AP Hand AR Whole AP Whole AR
Specalist SimpleBase [146] ResNet-50 65.2 73.8 61.5 74.9 60.6 71.5 46.0 58.4 52.1 63.3
SimpleBase [146] ResNet-152 68.2 76.4 66.1 78.7 62.3 72.8 48.1 60.7 54.8 66.1
HRNet [43] HRNet-W32 67.8 75.5 54.3 66.1 63.0 70.8 46.7 56.6 53.6 63.6
HRNet [43] HRNet-W48 70.1 77.6 67.5 78.7 65.6 74.3 53.5 63.9 57.9 68.1
TCFormer [147] TCFormer 69.1 77.0 69.8 81.3 64.9 74.6 53.5 65.0 57.2 67.8
RTMPose-m [148] CSPNeXt-m 67.3 75.0 61.5 75.2 81.3 87.1 47.5 58.9 58.2 67.4
DWPose-m [45] CSPNeXt-m 68.5 76.1 63.6 77.2 82.8 88.1 52.7 63.4 60.6 69.5
ViTPose-B [127] ViT-B 69.6 - 70.1 - 62.1 - 50.8 - 56.3 -
Generalist Hulk-FT ViT-B 70.2 77.5 65.9 79.1 84.2 90.1 53.2 63.1 61.6 72.1
D-B ICFG-PEDES
To further evaluate the effectiveness of Hulk on the pedestrian captioning task, we conduct transfer learning with ICFG-PEDES [149], another pedestrian dataset with detailed language captions. All pretrained image caption methods are fine-tuned on the training set with five epochs, and evaluated on the test set. As shown in Table XIX, Hulk outperforms powerful BLIP [141] and BLIP2 [79] with 3.2 B@4 and 1.5 B@4, 20.1 CIDEr and 12.8 CIDEr, respectively. Using fewer vision-language pretraining data (12M vs. 129M), the superior transfer performance of Hulk demonstrates its general capabilities on human-centric perception tasks.
TABLE XIX: Pedestrian image caption finetuning results on ICFG-PEDES [149]. * denotes fixed backbone.
Method Backbone B@4 CIDEr Trainable Params Total Params
Specialist BLIP [141] ViT-B 22.5 36.9 446.0M 446.0M
BLIP-2 [79] ViT-G* 24.2 44.2 188.0M 1.1B
Generalist Hulk ViT-B 25.7 57.0 113.7M 113.7M
Appendix E Model Parameters and FLOPs
We compute the model parameters and FLOPs of Hulk and some representative methods on all eight human-centric tasks in Table XX. Hulk-B has 113.7M parameters which is comparable with SOTA methods on most human-centric tasks. The only exceptions are skeleton-based action recognition and pedestrian image caption tasks. Compared with specialist models for the skeleton-based action recognition task, Hulk needs more parameters to enable its general abilities on other more complex human-centric tasks. Compared with specialist models for the pedestrian image caption task, notably, Hulk achieves comparable performance but with a more efficient architecture that eliminates the need for an additional text encoder, resulting in fewer parameters and lower computational requirements(GFLOPs).
TABLE $X$: Model parameters and GFLOPs of Hulk and other methods. GFLOPs are computed with the input size reported in corresponding papers. †denotes fixed backbone. ‡We compute the GFLOPs of autoregressively forwarding one step in the pedestrian image caption task.
Task Dataset Method Backbone Performance Params GFLOPs
Detection CrowdHuman DETR ResNet-50 75.9% mAP 41.0M 187.0
DDETR ResNet-50 91.5% mAP 40.0M 173.0
PATH-B ViT-B 90.9% mAP 106.0M 341.3
Hulk-B ViT-B 92.4% mAP 113.7M 348.2
2D Pose COCO HRNet HRNET-W48 75.1% AP 63.6M 32.9
TokenPose-L/D24 HRNET-W48-S3 75.8% AP 27.5M 11.0
ViTPose-B ViT-B 75.8% AP 86.0M 17.1
PATH-B ViT-B 76.3% AP 106.0M 18.2
Hulk-B ViT-B 77.5% AP 113.7M 19.1
Parsing LIP SCHP ResNet-101 59.4% mIoU 66.7M 77.2
PCNet ResNet-101 57.0% mIoU 53.8M 232.5
PATH-B ViT-B 61.4% mIoU 106.0M 87.3
Hulk-B ViT-B 64.0% mIoU 113.7M 91.4
Attribute PA-100K C-Tran ResNet-101 81.5% mA 67.8M 11.1
DAFL ResNet-50 83.5% mA 154.2M 24.8
PATH-B ViT-B 86.9% mA 106.0M 18.2
Hulk-B ViT-B 87.9% mA 113.7M 19.1
Skeleton NTU60-XSub ST-GCN GCN 81.5% acc 1.2M 1.7
AGCN GCN 88.5% acc 1.6M 2.1
PoseConv3D CNN 93.1% acc 2.0M 15.9
Hulk-B ViT-B 94.0% acc 113.7M 20.1
3D Pose & Mesh Human3.6M METRO HRNet-W64 56.5 MPJPE 230.4M 27.4
MeshGraphormer HRNet-W64 51.2 MPJPE 226.5M 28.9
FastMETRO HRNet-W64 52.2 MPJPE 153.0M 21.2
Hulk-B ViT-B 43.6 MPJPE 113.7M 19.7
Caption CUHK-PEDES BLIP ViT-B 32.9 B@4 446.0M 506.1‡
BLIP-2 ViT-G† 32.8 B@4 1.1B 2814.7‡
Hulk-B ViT-B 31.1 B@4 113.7M 58.7‡
Appendix F Additional Visualization
F-A Visualization of Eight Tasks
In Figure 10, we utilize Hulk to simultaneously predict eight human-centric tasks on the same input image, demonstrating excellent performance across all tasks simultaneously.
Figure 10: Visualization of Hulk’s performance on eight human-centric tasks for the same input.
F-B Visualization For Task Collaboration
To qualitatively demonstrate how one task helps the other in Hulk, we provide visualizations in two ways, i.e., visualization of prediction results and intermediate feature maps in diverse tasks.
F-B1 Visualization of prediction results in different tasks
We compare the predicted results on small tasks, e.g., pedestrian detection and attribute recognition tasks, with different training sets to visualize potential task collaborations: a) one dataset from a small task; b) one dataset from a small task + one dataset from a generic task; and c) datasets from 8 tasks. In the small ablation training set of Hulk, we choose detection and attribute recognition as small tasks and choose 2D pose estimation and 3D pose estimation as generic tasks:
  • We provide qualitative visualizations for pedestrian detection on the CrowdHuman validation set to visualize potential collaborations of tasks. As shown in Figure 11, the dashed boxes indicate missed detection. It can be observed that when trained on the single detection task, the model fails to detect some occluded individuals or small individuals (column 5). However, Hulk can detect previously missed individuals by jointly training additional generic tasks (such as 2D pose and 3D pose) and pedestrian detection tasks (columns 3 and 4). Furthermore, when Hulk is jointly trained with all eight tasks, all previously missed individuals are successfully detected (column 2), indicating potential collaborations among human-centric tasks.
  • We also provide the qualitative visualizations for another smaller task, i.e., pedestrian attribution recognition. As shown in Figure 12 (column 5), when only training on the attribute recognition task, Hulk misses some easy attributes (marked blue) and predicts wrong attributes (marked red) in the images, e.g., boots, backpacks, and shoulder bags. However, when 2D or 3D pose estimation tasks are incorporated into the joint training, most of these issues are resolved (columns 3 and 4). Furthermore, when we train Hulk on all 8 tasks (see column 2), even highly concealed attributes can be recognized, such as the UpperStride (stride T-shirt, a specific attribute in PA-100K) that the man wears under the coat in the second-row. These visualizations indicate that jointly training with generic tasks can benefit the learning on pedestrian attribute recognition, which is also a small task.
F-B2 Visualization of intermediate feature maps
In Figure 13, we visualize the feature map of the shared decoder’s last layer to confirm the consistency of regions of interest across different human-centric tasks. The similarities of intermediate feature maps help to visualize the improved performance of small tasks with the help of some generic tasks.
Specifically, 2D pose estimation and 3D pose estimation tasks highlight similar regions around human body joints, illustrating the high correlation between these two tasks. The phenomenon is consistent with the quantitative results in Table XII, where removing the 2D pose estimation task during joint training leads to a notable increased error (+16.7 MPJPE) in 3D pose estimation. Additionally, the feature map of 2D pose estimation task shows clear boundaries of the pedestrians (the low light part between the background and pedestrians), which helps us to understand why removing 2D pose task significantly decreases performance (-16.8% AP) in detection in Table XII. On the contrary, the feature maps of attribute and mesh recovery tasks show a low similarity. We suggest that features of mesh recovery can not help the model highlight regions of attributes, e.g., the upper clothing of the man in see the top example in Figure 13, resulting in slight performance changes (+0.1 mA) in attribute recognition after removing the mesh recovery task.
Figure 11: Visualization of pedestrian detection tasks on CrowdHuman [102] dataset. The dashed boxes indicate missed detections.
Figure 12: Visualization of pedestrian attribute recognition task on the PA-100K [101] dataset. The missed and incorrectly identified attributes are marked in blue and red, respectively.
Figure 13: Visualization of the feature map from the last layer of the decoder on seven human-centric tasks.
Appendix G Details of training datasets
G-A Full setting
As shown in Table XXI, we utilize 42 publically available datasets to form the training set for training our Hulk model, containing 30,187,836 training samples and covering eight different human-centric tasks, including human parsing, 2D pose estimation, attribute recognition, pedestrian detection, skeleton-based action recognition, image caption, 3D pose estimation, and mesh recovery.
G-B Ablation setting
In ablation studies, to demonstrate the effectiveness of Hulk and further explore the potential of a human-centric foundation model, we conduct several ablations on a smaller training set. In Table XXII, the ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks.
TABLE XXI: Statistics of Training Datasets in the full setting.